# Formal Degrees and Parabolic Induction: the Maximal Generic Case

### v2

@@ 4.2. Heiermann’s formula
In [Hei04, §8.6], as a byproduct of his spectral decomposition, Heiermann established a formula of $\deg(\pi)/\deg(\sigma)$ in terms of the residues of the $\mu$-function. To apply his formula to our case, we need some preparations.
First let $W=W^{\mathbf{G}}(\mathbf{T}_{0})$ be the Weyl group of $(\mathbf{G},\mathbf{T}_{0})$. Following notations of [Hei04], let
@@
  • $X^{\mathrm{ur}}(M)^{G}$ (resp. $X_{0}^{\mathrm{ur}}(M)^{G}$) be the image of $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}$ (resp. $i\mathfrak{a}_{\mathbf{M}}^{\mathbf{G}*}$) under the map $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}\twoheadrightarrow X^{\mathrm{ur}}(M)$ (note that the group $X^{\mathrm{ur}}(M)^{G}$ here is denoted by $X_{M}^{G}$ in [MW95, pp. 7]);
@@
as introduced in [Hei04, §6.1]. Since $|W(\mathbf{M})|\leq 2$, and by Shahidi’s classification $w_{0}\sigma\simeq\sigma$, we get the following
@@
Next we let $y_{0}\in\mathbb{R}^{\times}_{>0}$ be the smallest positive real number such that $\sigma\otimes\chi_{iy_{0}\tilde{\alpha}}\simeq\sigma$, so that $y_{0}\in\mathbb{Q}^{\times}_{>0}\cdot\left(\frac{2\pi}{\log q}\right)$ and
$\mathcal{O}_{0}^{G}=\{\sigma\otimes\chi_{iy_{0}\tilde{\alpha}}\mid y\in[0,y_{0})\}.$
Fixing the basis $\tilde{\alpha}$ of $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}=\mathbb{C}\cdot\tilde{\alpha}$. By a result of Silberger [Sil80, pp. 572], $\tilde{\alpha}/j$ is the only pole of the $\mu$-function on $\mathbb{R}_{>0}^{\times}\tilde{\alpha}\subset\mathfrak{a}_{\mathbf{M}}^{\mathbf{G}*}$ (where we identify $\mu$ as a rational function on $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}$). Now applying the residue theorem in complex analysis, we have
$\int_{0}^{y_{0}}\mu(\sigma\otimes\chi_{(r+iy)\tilde{\alpha}})dy=\int_{0}^{y_{0}}\mu(\sigma\otimes\chi_{iy\tilde{\alpha}})dy+2\pi\cdot\underset{s=1/j}{\mathrm{Res}}\mu(\sigma\otimes\chi_{s\tilde{\alpha}}).\quad(r\gg 0)$
@@
$\int_{\mathcal{O}_{0}^{G}}\mu(\sigma\otimes\chi_{r\tilde{\alpha}})d\sigma=\int_{\mathcal{O}_{0}^{G}}\mu(\sigma)d\sigma+\frac{\mathrm{vol}(\mathcal{O}_{0}^{G})}{y_{0}}\cdot 2\pi\cdot\underset{s=1/j}{\mathrm{Res}}\mu(\sigma\otimes\chi_{s\tilde{\alpha}}).\quad(r\gg 0)$
In order to make explicit use of Heiermann’s formula, we shall next compute the quotient $\mathrm{vol}(\mathcal{O}_{0}^{G})/y_{0}$ explicitly.
@@
This is a finite subgroup of $X_{0}^{\mathrm{ur}}(M)^{G}$. We denote its order by $t$. Then the smallest positive number $s_{0}\in\mathbb{R}_{>0}^{\times}$ such that
$\sigma\otimes|\chi|^{is_{0}}\simeq\sigma$
@@
$s_{0}=\left(\frac{2\pi}{\log q}\right)\cdot\frac{1}{lt}.$
Since $\tilde{\alpha}=\langle\chi,\alpha^{\vee}\rangle^{-1}\chi$, the smallest $y_{0}\in\mathbb{R}^{\times}_{>0}$ such that
$\sigma\otimes\chi_{iy_{0}\tilde{\alpha}}\simeq\sigma$
(recall that $\chi_{iy_{0}\tilde{\alpha}}$ is the image of $iy_{0}\tilde{\alpha}$ under $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}\twoheadrightarrow X_{0}^{\mathrm{ur}}(M)^{G}$) is then
$y_{0}=\langle\chi,\alpha^{\vee}\rangle\left(\frac{2\pi}{\log q}\right)\cdot\frac{1}{lt}.$
@@
$\mathrm{vol}(X_{0}^{\mathrm{ur}}(M)^{G})=\frac{m}{k},~{}\mathrm{vol}(\mathcal{O}_{0}^{G})=\frac{m}{lt}.$
@@
$\frac{\mathrm{vol}(\mathcal{O}_{0}^{G})}{y_{0}}=\left(\frac{\log q}{2\pi}\right)\cdot\frac{m}{\langle\chi,\alpha^{\vee}\rangle}.$
Recall that we’ve fixed a special maximal compact subgroup $K_{G}$ of $G$ at the beginning of this section, and for any closed subgroup $H$ of $G$, $H$ is equipped with the measure such that $K_{H}:=H\cap K_{G}$ has volume one. We denote by $\deg(\cdot)$ the formal degree with respect to this choice of the measure. Then by Heiermann’s formula in [Hei04, §8.6], we obtain the following
@@
$\frac{\deg(\pi)}{\deg(\sigma)}$ $=\gamma(\mathbf{G}/\mathbf{M})\cdot|\mathrm{Stab}(A)|^{-1}\cdot\frac{\mathrm{vol}(\mathcal{O}_{0}^{G})}{y_{0}}\cdot 2\pi\cdot\underset{s=1/j}{\mathrm{Res}}\mu(\sigma\otimes\chi_{s\tilde{\alpha}})$
@@ 5.1. Conditional Construction of Discrete Parameters
Assuming the LLC for $\mathbf{G}$ (resp. $\mathbf{M}$), which is conjectural in general, let $\varphi_{\pi}:\mathrm{WD}_{F}\rightarrow{{}^{L}G}$ (resp. $\varphi_{\sigma}:\mathrm{WD}_{F}\rightarrow{{}^{L}\!M}$) be the L-parameter of $\pi$ (resp. $\sigma$). Let ${{}^{L}\!P}$ be a parabolic subgroup of ${{}^{L}G}$ corresponding to $\mathbf{P}$ with inclusion $\iota_{P}:{{}^{L}\!P}\hookrightarrow{{}^{L}G}$; these are unique up to $\hat{G}$-conjugacy. The conjectural LLC includes the following properties:
@@
When $\pi$ is a discrete series and $\sigma$ is unitary supercuspidal, it is a natural question to ask how to construct $\varphi_{\pi}$ from $\varphi_{\sigma}$. This problem is much subtler than the above two standard cases, which was studied by Heiermann in [Hei06][Hei06a] using tools from local harmonic analysis developed by himeself. We now summarize his results as follows.
@@
Let $T_{{}^{L}\!M}$ be the maximal torus of $Z(\hat{M})^{\Gamma}$. Let $s_{\lambda}\in T_{{}^{L}\!M}$ be the element corresponding to $(-\lambda)\in\mathfrak{a}_{\mathbf{M}}^{*}$ by the LLC for tori (cf. §2.3, especially the final remark), and
@@
Let $\hat{M}^{\sigma}:=Z_{\hat{G}}(\varphi_{\sigma}(\mathrm{W}_{F}))$; this is a reductive Lie group (cf. [Hei06, Proposition 5.2]) but not necessarily connected in general. We denote by $\hat{M}^{\sigma}_{\mathrm{der}}$ its derived subgroup.
Under some natural assumptions (cf. [Hei06, §5.2, 5.3]) on irreducible supercuspidal representations of Levi subgroups, Heiermann proved that there exists a nilpotent element $N_{\sigma,\lambda}\in\mathrm{Lie}(\hat{M}^{\sigma}_{\mathrm{der}})$ (not unique in general), such that
@@
    a torus in ${{}^{L}G}$ that centralizes $s_{\sigma,\lambda}$ and $N_{\sigma,\lambda}$ simultaneouly is contained in $Z(\hat{M}^{\sigma}_{\mathrm{der}})$,
i.e. $(s_{\sigma,\lambda},N_{\sigma,\lambda})$ is an $L^{2}$-pair (for more details of $L^{2}$-pair, cf. [Hei06, §3.1][Lus83]). It was proved that such an $L^{2}$-pair determines an algebraic homomorphism $\psi_{\sigma,\lambda}:\mathrm{SL}_{2}(\mathbb{C})\rightarrow\hat{M}^{\sigma}$, with
@@
Theorem 5.1 (Heiermann, [Hei06]).
Under a series of natural hypotheses of the LLC and L-packets for $\mathbf{G}$ and its semi-standard Levi subgroups (for precise statements, cf. [Hei06, §6.5]), the map $\mathrm{LL}^{\mathbf{G}}$ constructed above does induce a bijection between the set of L-packets of $\mathbf{G}$ and $\Phi(\mathbf{G})$.
@@
The main input of the proof of this theorem is a characterization of the cuspidal support of discrete series via harmonic analysis. It claims that an irreducible supercuspidal of $M$ is the cuspidal support of some discrete series $\pi$ of $G$ if and only if the $\mu$-function has a pole of maximal order at $\sigma$ (cf. [Hei04]).
For generic representations of a quasi-split group, the compatibility of certain local factors associated to the above construction (namely $L(s,r_{i}\circ\varphi_{\pi})$, $\varepsilon(s,r_{i}\circ\varphi_{\pi},\psi)$, $\gamma(s,r_{i}\circ\varphi_{\pi},\psi)$ where $r_{i}$ as defined in §4.1) with Langlands-Shahidi local factors was studied in [Hei06a].
This result to some extent characterizes what a “reasonable” construction of discrete parameters (from supercuspidal ones) “should” be, while the precise statements of those hypotheses in the theorem are quite long and technical. However, we shall not make any use of them in the sequel (thus omitted here).
@@ 5.2. The Adjoint $\gamma$-Factors
  • $\pi$ a discrete series representation of $G$ which is a subrepresentation of $i_{P}^{G}\sigma_{\lambda}$ for $\lambda=s_{0}\tilde{\alpha}$, cf. §4.1;
@@
Of course, the LLC is inaccessible in general, so we need some “working hypotheses” in such a general setting. It is a well-known folklore conjecture that generic supercuspidal representations should correspond to $\mathrm{SL}_{2}$-trivial discrete parameters (cf. [Hei06a, pp. 103] or more formally, Conjecture 7.1 of [GR10]). Thus we assume further that
@@
For a proof, see [Hei06a, §3.2], and [Sha90, §3.4]. In particular, we have the following
@@
Now let ${{}^{L}\mathfrak{g}}$ be the Lie algebra of $\hat{G}$ and $\mathrm{Ad}$ be the adjoint representation of ${}^{L}G$ on ${{}^{L}\mathfrak{g}}_{\mathrm{ad}}:={{}^{L}\mathfrak{g}}/Z({{}^{L}\mathfrak{g}})^{\Gamma}$. We consider the composite
$\mathrm{W}_{F}\xrightarrow[]{\varphi_{\pi}^{\prime}}{{}^{L}G}\xrightarrow{\mathrm{Ad}}\mathrm{GL}({{}^{L}\mathfrak{g}}_{\mathrm{ad}}).$
@@
we first decompose the restriction of $\mathrm{Ad}$ onto ${{}^{L}\!M}$. Since ${{}^{L}\mathfrak{g}}={{}^{L}\mathfrak{m}}\oplus{{}^{L}\mathfrak{n}}\oplus{{}^{L}\overline{\mathfrak{n}}}$ and $Z({{}^{L}\mathfrak{g}})^{\Gamma}\subset Z({{}^{L}\mathfrak{m}})^{\Gamma}$, the adjoint action of ${{}^{L}\!M}$ on ${{}^{L}\mathfrak{g}}_{\mathrm{ad}}$ can be decomposed as
${{}^{L}\mathfrak{g}}_{\mathrm{ad}}={{}^{L}\mathfrak{m}}/Z({{}^{L}\mathfrak{g}})^{\Gamma}\oplus{{}^{L}\mathfrak{n}}\oplus{{}^{L}\overline{\mathfrak{n}}}$
as an ${}^{L}\!M$-module. Since $\mathbf{P}=\mathbf{M}\mathbf{N}$ is a standard maximal parabolic subgroup, ${{}^{L}\!P}$ is relevant, maximal, $Z({{}^{L}\mathfrak{m}})^{\Gamma}/Z({{}^{L}\mathfrak{g}})^{\Gamma}$ is then one dimensional and isomorphic to the trivial representation $\mathbbm{1}$ of ${}^{L}\!M$.
Recall by the Langlands-Shahidi theory, the adjoint representation of ${}^{L}\!M$ on ${{}^{L}\mathfrak{n}}$, denoted by $r$, can be decomposed as $(r,{{}^{L}\mathfrak{n}})=\bigoplus_{i=1}^{m}(r_{i},V_{i})$, where
$V_{i}:=\{X_{\beta^{\vee}}\in{{}^{L}\mathfrak{n}}\mid\beta\in\Sigma,\langle\tilde{\alpha},\beta^{\vee}\rangle=i\},\quad 1\leq i\leq m,$
and each $(r_{i},V_{i})$ is irreducible. Similarly, $(\tilde{r},{{}^{L}\overline{\mathfrak{n}}})=\bigoplus_{i=1}^{m}(\tilde{r}_{i},\overline{V}_{i})$, where $\tilde{r}$ (resp. $\tilde{r}_{i}$) is the contragredient of $r$ (resp. $r_{i}$), and
$\overline{V}_{i}:=\{X_{\beta^{\vee}}\in{{}^{L}\overline{\mathfrak{n}}}\mid\beta\in\Sigma,\langle\tilde{\alpha},\beta^{\vee}\rangle=(-i)\},\quad 1\leq i\leq m.$
@@
where we denote by $\mathrm{Ad}_{M}$ the adjoint representation of ${{}^{L}\!M}$ on ${{}^{L}\mathfrak{m}}/Z({{}^{L}\mathfrak{m}})^{\Gamma}$. By definition of $s_{\lambda}\in T_{{}^{L}\!M}$, it corresponds to $\lambda=-s_{0}\tilde{\alpha}\in\mathfrak{a}_{\mathbf{M}}^{*}$ by the LLC for tori (cf. §2.3). By our assumption
@@
is trivial, thus $s_{\sigma,\lambda}:=s_{\lambda}s_{\sigma}=s_{\lambda}\in Z({{}^{L}\!M})$ acts trivially on ${{}^{L}\mathfrak{m}}$, $\mathrm{Ad}_{M}\circ\varphi_{\pi}^{\prime}\simeq\mathrm{Ad}_{M}\circ\varphi_{\sigma}$ and
@@
$V_{i}=\{X_{\beta^{\vee}}\in{{}^{L}\mathfrak{n}}\mid\beta\in\Sigma,\langle\tilde{\alpha},\beta^{\vee}\rangle=i\},$
@@
As explained in Lemma 1.2 of [HII08a], $\gamma(0,\pi,\mathrm{Ad},\psi),\gamma(0,\sigma,\mathrm{Ad},\psi)$ are holomorphic and nonzero at $s=0$. Since when $s\rightarrow 0$

### v3

@@ 4.2. Heiermann’s formula
In [heiermann-spectrales, §8.6], as a byproduct of his spectral decomposition, Heiermann established a formula of $\deg(\pi)/\deg(\sigma)$ in terms of the residues of the $\mu$-function. To apply his formula to our case, we need some preparations.
First let $W=W^{\mathbf{G}}(\mathbf{T}_{0})$ be the Weyl group of $(\mathbf{G},\mathbf{T}_{0})$. Following notations of [heiermann-spectrales], let
@@
  • $X^{\mathrm{ur}}(M)^{G}$ (resp. $X_{0}^{\mathrm{ur}}(M)^{G}$) be the image of $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}$ (resp. $i\mathfrak{a}_{\mathbf{M}}^{\mathbf{G}*}$) under the map $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}\twoheadrightarrow X^{\mathrm{ur}}(M)$ (note that the group $X^{\mathrm{ur}}(M)^{G}$ here is denoted by $X_{M}^{G}$ in [eisenstein, pp. 7]);
@@
as introduced in [heiermann-spectrales, §6.1]. Since $|W(\mathbf{M})|\leq 2$, and by Shahidi’s classification $w_{0}\sigma\simeq\sigma$, we get the following
@@
Next we let $y_{1}\in\mathbb{R}^{\times}_{>0}$ be the smallest positive real number such that $\sigma\otimes\chi_{iy_{1}\tilde{\alpha}}\simeq\sigma$, so that $y_{1}\in\mathbb{Q}^{\times}_{>0}\cdot\left(\frac{2\pi}{\log q}\right)$ and
$\mathcal{O}_{0}^{G}=\{\sigma\otimes\chi_{iy_{1}\tilde{\alpha}}\mid y\in[0,y_{1})\}.$
We fix the basis $\tilde{\alpha}$ of $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}=\mathbb{C}\cdot\tilde{\alpha}$.
In our case, the induced representation $i_{P}^{G}\sigma_{\tilde{\alpha}/s}$ is irreducible for $s>0$ and $s\neq 1/j$ by Theorem 8.1 of [shahidi-plancherel]. Hence on the half line $s\tilde{\alpha}$, $s>0$, the unique pole of $\mu(\sigma\otimes\chi_{s\tilde{\alpha}})$ is $s=1/j$ ([silberger, Lemma 5.4.2.4]). Now applying the residue theorem in complex analysis to the contour
[…]
we get
$\int_{0}^{y_{1}}\mu(\sigma\otimes\chi_{(r+iy)\tilde{\alpha}})dy=\int_{0}^{y_{1}}\mu(\sigma\otimes\chi_{iy\tilde{\alpha}})dy+2\pi\cdot\underset{s=1/j}{\mathrm{Res}}\mu(\sigma\otimes\chi_{s\tilde{\alpha}}).\quad(r\gg 0)$
@@
(4.1) $\int_{\mathcal{O}_{0}^{G}}\mu(\sigma\otimes\chi_{r\tilde{\alpha}})d\sigma=\int_{\mathcal{O}_{0}^{G}}\mu(\sigma)d\sigma+\frac{\mathrm{vol}(\mathcal{O}_{0}^{G})}{y_{1}}\cdot 2\pi\cdot\underset{s=1/j}{\mathrm{Res}}\mu(\sigma\otimes\chi_{s\tilde{\alpha}}).\quad(r\gg 0)$
In order to make explicit use of Heiermann’s formula, we shall next compute the quotient $\mathrm{vol}(\mathcal{O}_{0}^{G})/y_{1}$ explicitly.
@@
This is a finite subgroup of $X_{0}^{\mathrm{ur}}(M)^{G}$. We denote its order by $t$ (the “torsion number”). Then the smallest positive number $s_{1}\in\mathbb{R}_{>0}^{\times}$ such that
$\sigma\otimes|\chi|^{is_{1}}\simeq\sigma$
@@
$s_{1}=\left(\frac{2\pi}{\log q}\right)\cdot\frac{1}{lt}.$
Since $\tilde{\alpha}=\langle\chi,\alpha^{\vee}\rangle^{-1}\chi$, the smallest $y_{1}\in\mathbb{R}^{\times}_{>0}$ such that
$\sigma\otimes\chi_{iy_{1}\tilde{\alpha}}\simeq\sigma$
(recall that $\chi_{iy_{1}\tilde{\alpha}}$ is the image of $iy_{1}\tilde{\alpha}$ under $\mathfrak{a}_{\mathbf{M},\mathbb{C}}^{\mathbf{G}*}\twoheadrightarrow X_{0}^{\mathrm{ur}}(M)^{G}$) is then
$y_{1}=\langle\chi,\alpha^{\vee}\rangle\left(\frac{2\pi}{\log q}\right)\cdot\frac{1}{lt}.$
@@
$\mathrm{vol}(X_{0}^{\mathrm{ur}}(M)^{G})=\frac{m}{l},~{}\mathrm{vol}(\mathcal{O}_{0}^{G})=\frac{m}{lt}.$
@@
$\frac{\mathrm{vol}(\mathcal{O}_{0}^{G})}{y_{1}}=\left(\frac{\log q}{2\pi}\right)\cdot\frac{m}{\langle\chi,\alpha^{\vee}\rangle}.$
Recall that we’ve fixed a special maximal compact subgroup $K_{G}$ of $G$ at the beginning of this section, and for any closed subgroup $H$ of $G$, $H$ is equipped with the measure such that $K_{H}:=H\cap K_{G}$ has volume one. We denote by $\deg(\cdot)$ the formal degree with respect to this choice of the measure. Then by Heiermann’s formula in [heiermann-spectrales, §8.6], we obtain the following
@@
$\frac{\deg(\pi)}{\deg(\sigma)}$ $=\gamma(\mathbf{G}/\mathbf{M})\cdot|\mathrm{Stab}(A)|^{-1}\cdot\frac{\mathrm{vol}(\mathcal{O}_{0}^{G})}{y_{1}}\cdot 2\pi\cdot\underset{s=1/j}{\mathrm{Res}}\mu(\sigma\otimes\chi_{s\tilde{\alpha}})$
@@
Proof.
The general formula of [heiermann-spectrales, §8.6] reads
$\frac{\deg(\pi)}{\deg(\sigma)}=\gamma(\mathbf{G}/\mathbf{M})\cdot|\mathrm{Stab}(A)|^{-1}\cdot m(\pi,\sigma)\cdot(\mathrm{Res}_{A}^{P}\mu)(\sigma).$
The multiplicity $m(\pi,\sigma)$ above in our case is one, since $\pi$ is the unique generic subquotient of $i_{P}^{G}\sigma_{\tilde{\alpha}/j}$ (cf. (3) of Theorem 4.1). The only point that requires an explanation is the notation $(\mathrm{Res}_{A}^{P}\mu)(\sigma)$ of Heiermann.
This is a multiple residue operator defined in a more general setting for arbitrary $\mathbf{P}$, not necessarily maximal. The general definition is quite tough, but in our case, since $\mathbf{P}$ is maximal (i.e. of corank one), up to constant this operator is exactly $\mathrm{Res}_{s=1/j}\mu(\sigma\otimes\chi_{s\tilde{\alpha}})$, and by [heiermann-spectrales, §4.11] (we take $\psi$ there to be $\mu$, and as we mentioned before, $\tilde{\alpha}/j\in\mathbb{C}\cdot\tilde{\alpha}$ is the unique pole of $\mu$ on the right half plane) it satisfies
$\int_{\mathcal{O}_{0}^{G}}\mu(\sigma\otimes\chi_{r\tilde{\alpha}})d\sigma=\int_{\mathcal{O}_{0}^{G}}\mu(\sigma)d\sigma+(\mathrm{Res}_{A}^{P}\mu)(\sigma).\quad(r\gg 0)$
Thus by our computation (4.1) above, it is exactly
$(\mathrm{Res}_{A}^{P}\mu)(\sigma)=\frac{\mathrm{vol}(\mathcal{O}_{0}^{G})}{y_{1}}\cdot 2\pi\cdot\underset{s=1/j}{\mathrm{Res}}\mu(\sigma\otimes\chi_{s\tilde{\alpha}}),$
hence the result. ∎
We remark that the notation “$\tilde{\alpha}$” introduced in [heiermann-spectrales, §3.2] is not the fundamental weight (following the convention of Shahidi) corresponding to $\alpha$ in our case. Since it is used to define the multiple residue operator in general, one has to be careful while applying the formula to more general cases.
@@ 5.1. Conditional Construction of Discrete Parameters
Assuming the LLC for $\mathbf{G}$ (resp. $\mathbf{M}$), which is conjectural in general, let $\varphi_{\pi}:\mathrm{WD}_{F}\rightarrow{{}^{L}G}$ (resp. $\varphi_{\sigma}:\mathrm{WD}_{F}\rightarrow{{}^{L}\!M}$) be the L-parameter of $\pi$ (resp. $\sigma$). Let ${{}^{L}\!P}$ be a (relevant) parabolic subgroup of ${{}^{L}G}$ corresponding to $\mathbf{P}$ with inclusion $\iota_{P}:{{}^{L}\!P}\hookrightarrow{{}^{L}G}$; these are unique up to $\widehat{G}$-conjugacy. The conjectural LLC includes the following properties:
@@
When $\pi$ is a discrete series and $\sigma$ is unitary supercuspidal, it is a natural question to ask how to construct $\varphi_{\pi}$ from $\varphi_{\sigma}$. This problem is much subtler than the above two standard cases, which was studied by Heiermann in [heiermann-orbites][heiermann-unipotent] using tools from local harmonic analysis developed by himeself. We now summarize his results as follows.
@@
Let $T_{{}^{L}\!M}$ be the maximal torus of $Z(\widehat{M})^{\Gamma}$. Let $s_{\lambda}\in T_{{}^{L}\!M}$ be the element corresponding to $(-\lambda)\in\mathfrak{a}_{\mathbf{M}}^{*}$ by the LLC for tori (cf. §2.3, especially the final remark), and
@@
Let $\widehat{M}^{\sigma}:=Z_{\widehat{G}}(\varphi_{\sigma}(\mathrm{W}_{F}))$; this is a complex reductive Lie group (cf. [heiermann-orbites, Proposition 5.2]) but not necessarily connected in general. We denote by $\widehat{M}^{\sigma}_{\mathrm{der}}$ its derived subgroup.
Under some natural assumptions (cf. [heiermann-orbites, §5.2, 5.3]) on irreducible supercuspidal representations of Levi subgroups, Heiermann proved that there exists a nilpotent element $N_{\sigma,\lambda}\in\mathrm{Lie}(\widehat{M}^{\sigma}_{\mathrm{der}})$ (not unique in general), such that
@@
    a torus in ${{}^{L}G}$ that centralizes $s_{\sigma,\lambda}$ and $N_{\sigma,\lambda}$ simultaneouly is contained in $Z(\widehat{M}^{\sigma}_{\mathrm{der}})$,
i.e. $(s_{\sigma,\lambda},N_{\sigma,\lambda})$ is an $L^{2}$-pair (for more details of $L^{2}$-pair, cf. [heiermann-orbites, §3.1][lusztig]). It was proved that such an $L^{2}$-pair determines an algebraic homomorphism $\psi_{\sigma,\lambda}:\mathrm{SL}_{2}(\mathbb{C})\rightarrow\widehat{M}^{\sigma}$, with
@@
Theorem 5.1 (Heiermann, [heiermann-orbites]).
Under a series of natural hypotheses of the LLC and L-packets for $\mathbf{G}$ and its semi-standard Levi subgroups (for precise statements, cf. [heiermann-orbites, §6.5]), the map $\mathrm{LL}^{\mathbf{G}}$ constructed above does induce a bijection between the set of L-packets of $\mathbf{G}$ and $\Phi(\mathbf{G})$.
@@
The main input of the proof of this theorem is a characterization of the cuspidal support of discrete series via harmonic analysis. It claims that an irreducible supercuspidal representation of $M$ is the cuspidal support of some discrete series $\pi$ of $G$ if and only if the $\mu$-function has a pole of maximal order at $\sigma$ (cf. [heiermann-spectrales]).
For generic representations of a quasi-split group, the compatibility of certain local factors associated to the above construction (namely $L(s,r_{i}\circ\varphi_{\pi})$, $\varepsilon(s,r_{i}\circ\varphi_{\pi},\psi)$, $\gamma(s,r_{i}\circ\varphi_{\pi},\psi)$ where $r_{i}$ as defined in §4.1) with Langlands-Shahidi local factors was studied in [heiermann-unipotent].
This result to some extent characterizes what a “reasonable” construction of discrete parameters (from supercuspidal ones) “should” be, while the precise statements of those hypotheses in the theorem are quite long and technical. However, we shall not make any use of these precise statements in the sequel (thus omitted here).
@@ 5.2. The Adjoint $\gamma$-Factors
  • $\pi$ a discrete series representation of $G$ which is a subrepresentation of $i_{P}^{G}\sigma_{\lambda}$ for $\lambda=s_{0}\tilde{\alpha}$, $s_{0}\in\{1,1/2\}$ (cf. §4.1); here we write $s_{0}$ rather than $1/j$ to make the ideas and details below clearer (since the exact number $1/j$ is not crucial here);
@@
Of course, the LLC is inaccessible in general, so we need some “working hypotheses” in such a general setting. It is a well-known folklore conjecture that generic supercuspidal representations should correspond to $\mathrm{SL}_{2}$-trivial discrete parameters (cf. [heiermann-unipotent, pp. 103] or more formally, Conjecture 7.1 of [gross-reeder]). Thus we assume further that
@@
For a proof, see [heiermann-unipotent, §3.2], and [shahidi-plancherel, §3.4]. In particular, we have the following
@@
Now let $\widehat{\mathfrak{g}}$ be the Lie algebra of $\widehat{G}$ and $\mathrm{Ad}$ be the adjoint representation of ${}^{L}G$ on $\widehat{\mathfrak{g}}_{\mathrm{ad}}:=\widehat{\mathfrak{g}}/Z(\widehat{\mathfrak{g}})^{\Gamma}$ (not $\widehat{\mathfrak{g}}/Z(\widehat{\mathfrak{g}})$). We consider the composite
$\mathrm{W}_{F}\xrightarrow{\varphi_{\pi}^{\prime}}{{}^{L}G}\xrightarrow{\mathrm{Ad}}\mathrm{GL}(\widehat{\mathfrak{g}}_{\mathrm{ad}}).$
@@
we first decompose the restriction of $\mathrm{Ad}$ onto ${{}^{L}\!M}$. Since $\widehat{\mathfrak{g}}=\widehat{\mathfrak{m}}\oplus\widehat{\mathfrak{n}}\oplus\widehat{\overline{\mathfrak{n}}}$ and $Z(\widehat{\mathfrak{g}})^{\Gamma}\subset Z(\widehat{\mathfrak{m}})^{\Gamma}$, the adjoint action of ${{}^{L}\!M}$ on $\widehat{\mathfrak{g}}_{\mathrm{ad}}$ can be decomposed as
$\widehat{\mathfrak{g}}_{\mathrm{ad}}=\widehat{\mathfrak{m}}/Z(\widehat{\mathfrak{g}})^{\Gamma}\oplus\widehat{\mathfrak{n}}\oplus\widehat{\overline{\mathfrak{n}}}$
as an ${}^{L}\!M$-module. Since $\mathbf{P}=\mathbf{M}\mathbf{N}$ is a standard maximal parabolic subgroup, ${{}^{L}\!P}$ is relevant, maximal, $Z(\widehat{\mathfrak{m}})^{\Gamma}/Z(\widehat{\mathfrak{g}})^{\Gamma}$ is then one dimensional and isomorphic to the trivial representation $\mathbbm{1}$ of ${}^{L}\!M$.
Recall by the Langlands-Shahidi theory, the adjoint representation of ${}^{L}\!M$ on $\widehat{\mathfrak{n}}$, denoted by $r$, can be decomposed as $(r,\widehat{\mathfrak{n}})=\bigoplus_{i=1}^{m}(r_{i},V_{i})$, where
$V_{i}:=\{X_{\beta^{\vee}}\in\widehat{\mathfrak{n}}\mid\beta\in\Sigma,\langle\tilde{\alpha},\beta^{\vee}\rangle=i\},\quad 1\leq i\leq m,$
and each $(r_{i},V_{i})$ is irreducible. Similarly, $(\tilde{r},\widehat{\overline{\mathfrak{n}}})=\bigoplus_{i=1}^{m}(\tilde{r}_{i},\overline{V}_{i})$, where $\tilde{r}$ (resp. $\tilde{r}_{i}$) is the contragredient of $r$ (resp. $r_{i}$), and
$\overline{V}_{i}:=\{X_{\beta^{\vee}}\in\widehat{\overline{\mathfrak{n}}}\mid\beta\in\Sigma,\langle\tilde{\alpha},\beta^{\vee}\rangle=(-i)\},\quad 1\leq i\leq m.$
@@
where we denote by $\mathrm{Ad}_{M}$ the adjoint representation of ${{}^{L}\!M}$ on $\widehat{\mathfrak{m}}/Z(\widehat{\mathfrak{m}})^{\Gamma}$. By definition of $s_{\lambda}\in T_{{}^{L}\!M}$, it corresponds to $\lambda=-s_{0}\tilde{\alpha}\in\mathfrak{a}_{\mathbf{M}}^{*}$ by the LLC for tori (cf. §2.3). By our assumption
@@
is trivial, thus $s_{\sigma,\lambda}:=s_{\lambda}s_{\sigma}=s_{\lambda}\in Z({{}^{L}\!M})$ acts trivially on $\widehat{\mathfrak{m}}$, $\mathrm{Ad}_{M}\circ\varphi_{\pi}^{\prime}\simeq\mathrm{Ad}_{M}\circ\varphi_{\sigma}$ and
@@
$V_{i}=\{X_{\beta^{\vee}}\in\widehat{\mathfrak{n}}\mid\beta\in\Sigma,\langle\tilde{\alpha},\beta^{\vee}\rangle=i\},$
@@
As explained in Lemma 1.2 of [hiraga-ichino-ikeda], $\gamma(0,\pi,\mathrm{Ad},\psi),\gamma(0,\sigma,\mathrm{Ad},\psi)$ are holomorphic and nonzero at $s=0$. Since when $s\rightarrow 0$